# The Forest Behind the Tree: Revealing Hidden Smart Home Communication Patterns

### v2

@@ IV-D A closer look at DNS data
We focus on event signatures related to boot interactions, as those proved to provide the most information concerning DNS usage. Fig. 7 shows the data related to both questions: the count of unique domain names contacted, and that of DNS resolvers queried. In both cases, the first-level and the hidden count are shown. The results for all types of interactions is shown in Fig. 11 in the appendix.
@@ IV-D2 DNS resolvers
IV-E Comparison with related work
In this last part of the evaluation, we compare the profiles generated by our framework with those generated by methods proposed in related work, namely PingPong [11], BehavIoT [14], and MUDgee [42], for the devices present in both their testbed and ours. The former two contributions profile specific individual events, whereas the latter profiles the device as a whole. Table II summarizes the devices and interactions covered by each work.
@@ IV-E Comparison with related work
[42] [11] [14]
TP-Link HS110 [26] toggle toggle device
Philips Hue lamp [38] toggle device
SmartThings Outlet [32] toggle boot device
D-Link camera [37] stream
TABLE II: Related work device & interaction coverage
@@
Figure 8: Comparison of discovered unique Flow IDs between three related works and ours.
@@
Data from PingPong and MUDgee were taken from their reported results, whereas BehavIoT’s results were reproduced on our side. Fig. 8 compares the count of unique Flow IDs discovered by each work. The three covered related works, by design, do not explore the event signature tree deeper than the first level. Considering event-level profiling, our methodology is, as expected, able to extract more Flow IDs than PingPong and BehavIoT, thanks to a deeper tree inspection, and also a more thorough communication pattern extraction, even at the first level. For device-level profiling, the value representing our work has been computed by merging the event signature trees extracted from all events of the given device. For two of the instrumented devices, our strategy is again able to extract a significantly higher number of Flow IDs than MUDgee. However, the number of Flow IDs is identical for the TP-Link HS110. After manual investigation, it appears the extracted Flow IDs are different: 10 out of 14 MUDgee flows are NTP traffic towards various servers, whereas our solution did not extract any NTP patterns. This traffic might therefore be part of periodic patterns, which are not covered by our strategy. On the other hand, we extract 10 non-NTP traffic patterns which were not detected by MUDgee.
@@ V Related Work
In the following, we will summarize related works, and position them along the aforementioned abstraction axes; Table III shows the resulting classification.
@@
Flows This work MUD [6], Hamza [10], Mandalari [23] De Keersmaeker [9], Girish [4]
@@
TABLE III: Abstraction levels of device behavior models
@@
Regarding research efforts to extract event signatures from network traffic, three works follow a workflow similar to ours: PingPong [11], D-interact [43], and IoTa [44]. Nevertheless, all three lie at a finer granularity level than our work, as their signatures are composed of individual packets, and leverage packet-level metadata, such as the packet size. We argue that such features can be subject to network churn and therefore vary from one event execution to the other, making them imprecise to accurately characterize network behavior.
@@
All of the above-mentioned works have in common that, unlike our work, they do not actively attempt to discover hidden communication patterns.
@@ VI Conclusion
In this article, we presented a novel methodology and framework to extract multi-level signatures from Smart Home device events, whereas existing works are limited to first-level signatures. Our signatures take the form of trees, one per interaction type with the object, with nodes being the constitutive flows. We show that more than half of the devices exhibit hidden traffic patterns, only occurring when tampering with the network they are connected to. In average, two hidden flows are discovered for each device, across all studied interactions with the device. Our work highlights the robustness capabilities of Smart Home devices, enabling users to take more advised buying decisions. One can envisage that, in the future, devices are labeled with such a robustness score.
As future work, we would like to apply our methodology to scenarios where multiple Smart Home devices are involved in a single interaction. This is conceptually supported by our methodology, but will require an extension of the mechanisms we used for event triggering. In this case, approaches to further automate the latter should also be investigated.

### v3

@@ IV-D A closer look at DNS data
We focus on event signatures related to boot interactions, as those proved to provide the most information concerning DNS usage. Fig. 8 shows the data related to both questions: the count of unique domain names contacted, and that of DNS resolvers queried. In both cases, the first-level 222 In this context, the qualifier first-level indicates the occurrences at the signature tree’s first depth level; it has no relation with the similarly named, but unrelated, domain names’ top-level domain. and the hidden count are shown.
@@ IV-D2 DNS resolvers
IV-E Analysis of the node pruning heuristic
We analyze the effect of our tree pruning heuristic described in Section III-D2. Fig. 9 illustrates, for every studied event, the count of Flow IDs which have been pruned while generating the event signature tree, split per node depth. As a reminder, this heuristic states that we prune a node, i.e. we do not process the children of a node, if an equivalent node is already present in the tree; the depicted count is therefore equal to the number of duplicate Flow IDs in the tree.
Figure 9: Count of pruned Flow IDs, per depth and per event
We observe that the count is generally important. In terms of efficiency, this means that our heuristics manages to significantly reduce the time and resources taken by the event signature tree generation.
We also notice that most of the duplicate Flow IDs occur at depth 2. Those represent all the flows that occur after blocking a first-level flow. Their large number shows that most devices exhibit a high redundancy in communication patterns. Indeed, if the device must change its network behavior to circumvent a traffic restriction, it will usually only change the affected protocol, e.g. by contacting another domain name or switching from TCP to UDP; it will not completely modify the communication pattern.
IV-F Comparison with related work
We compare the profiles generated by our framework with those generated by methods proposed in related work, namely PingPong [11], BehavIoT [14], and MUDgee [43], for the devices present in both their testbeds and ours. PingPong and BehavIoT profile specific interactions of a device, whereas MUDgee profiles the device as a whole. Table II summarizes the devices and interactions covered by each work, as well as the smartphone apps which produced the subset of our data used for our side in the comparison.
@@ IV-F Comparison with related work
Devices Apps [43] [11] [14]
TL plug Kasa Smart toggle toggle device
Hue lamp Philips Hue toggle device
ST plug SmartThings toggle boot device
DL cam mydlink stream
TABLE II: Related works’ device and interaction coverage. Device legend: TL plug = TP-Link HS110; Hue lamp = Philips Hue lamp; ST plug = SmartThings Outlet; DL cam = D-Link camera.
@@
Figure 10: Comparison of discovered unique Flow IDs between three related works and ours. Note that MUDgee profiles devices as a whole and not for individual interactions.
@@
Fig. 10 compares the number of unique Flow IDs discovered by each work and by us. Data from PingPong and MUDgee were taken from their reported results, whereas BehavIoT’s results were reproduced on our side. The three related works, by design, do not explore the event signature tree deeper than the first level. Regarding the specific interactions, our approach is, as expected, able to extract more Flow IDs than PingPong and BehavIoT, thanks to a deeper tree inspection, and also a more thorough communication pattern extraction, even at the first level. To compare our results with the device-level profiles created by MUDgee, we count the unique Flow IDs over all interactions of a device. For two of the instrumented devices, namely the SmartThings Outlet and the Hue light, our approach is again able to extract a significantly higher number of Flow IDs than MUDgee. However, the number of Flow IDs is identical for the TP-Link HS110. A manual investigation shows that the extracted Flow IDs are different: 10 out of 14 MUDgee flows are NTP traffic towards various servers, whereas our solution did not extract any NTP patterns. This traffic might be part of periodic patterns, which are not covered by our approach; as explained in Section III-B, our methodology excludes traffic that does not appear in every iteration of a tested interaction from the profile building process. On the other hand, we extract 10 non-NTP traffic patterns which were not detected by MUDgee.
Additionally, we compare the number of domains names discovered by our approach with those reported by Mandalari et al. [24]. The latter focuses on extracting the hosts from Smart Home communication patterns. Table III shows, in the columns “Total ours” and “Total theirs”, the total number of unique domain names discovered by our approach and by theirs for the devices and interactions covered in both works. We observe that our multi-level approach is able to discover more domain names. However, it should be noted that changes in the firmware of the devices since the publication of the cited article [24] could influence the results.
Device & interact. Total ours Total theirs Both Ours only Theirs only
TP-Link HS110 boot & toggle 5 4 use1-api.tplinkra.com euw1-api.tplinkra.com euw1-device-telemetry-gw.iot.i.tplinknbu n-wap.tplinkcloud.com n-devs.tplinkcloud.com n-deventry.tplinkcloud.com
SmartThings Outlet boot 4 3 / hub.smartthings.com usher.connect.smartthings.com connectivity.smartthings.com dc2-eu01-euwest1.connect.smartthings.com fw-update2.smartthings.com
Philips Hue lamp boot 5 4 / mqtt-eu-01.iot.meethue.com time.meethue.com ntp2.aliyun.com ntp4.aliyun.com time4.google.com diagnostics.meethue.com ecdinterface.philips.com
TABLE III: Domain names extracted by Mandalari et al. [24] (referred to as “theirs”) and our work (referred to as “ours”). Note that the domain names reported by Mandalari et al. only concern non-essential hosts.
A direct comparison of the domain names beyond the mere numbers is unfortunately not possible. This is because Mandalari et al. were primarily interested in finding “non-essential” hosts, i.e. hosts that can be blocked, for example for privacy reasons, without negatively impacting the essential functionality of an IoT device. Consequently, while the total numbers are known, they only published the domain names of non-essential hosts, whereas we do not differ between essential and non-essential ones. Nevertheless, we have also included the actual domain names in Table III. The differences between the two works can be explained by the fact that our approach actively looks for hidden communication patterns, the fact that their work is focused on identifying non-essential hosts, and the reasons already mentioned above. For example, the domain name ecdinterface.philips.com does not exist anymore (in 2025), and the domain name diagnostics.meethue.com could be part of a periodic or sporadic pattern filtered out by our approach. This can also explain the empty intersections between the two works for two of the tested devices.
@@ V Related Work
In the following, we will summarize related works, and position them along the aforementioned abstraction axes; Table IV shows the resulting classification.
@@
Flows Mandalari [24], This work MUD [6], Hamza [10] De Keersmaeker [9], Girish [4]
@@
TABLE IV: Abstraction level of device behavior models in our work and in the literature
@@
Regarding research efforts to extract event signatures from network traffic, three works follow an interact&measure workflow similar to ours: PingPong [11], D-interact [44], and IoTa [45]. Nevertheless, all three lie at a finer granularity level than our work, as their signatures are composed of individual packets, and leverage packet-level metadata, such as the packet size. We argue that such features can be subject to network churn and therefore vary from one event execution to the other, making them imprecise to accurately characterize network behavior.
Mandalari et al. [24] centered their analysis on the hosts contacted by Smart Home devices, characterized mainly by their domain names, but falling back to their plain IP address if the domain name was unavailable. Their goal was to identify non-required hosts, i.e. hosts which could be blocked without impairing the device’s operation, e.g. telemetry or advertisement services. While they designed a workflow similar to ours to identify the successful events, they did not further explore the event signature tree.
@@
All the above-mentioned works have in common that, unlike our work, they do not actively attempt to discover hidden communication patterns.
@@ VI Conclusion
In this article, we presented a novel methodology and framework to extract comprehensive, multi-level traffic signatures from Smart Home device events, whereas existing works are limited to first-level signatures. The signatures take the form of trees, with the nodes being the constitutive traffic flows. By the usage of a dynamic blocking scheme, our approach allows uncovering hidden traffic patterns that only occur when the default communication mechanisms of a device fail.
We show that more than half of the tested devices exhibit such hidden traffic patterns. In average, two hidden flows are discovered for each device, across all studied interactions with the device. Our work highlights the robustness capabilities of Smart Home devices, by proposing a new metric dubbed the robustness score. We can envision that, in the future, such a score might be used as a marketing argument; indeed, customers would prefer a device which is able to function even in inhospitable network conditions.
As future work, we would like to apply our methodology to scenarios where multiple Smart Home devices are involved in a single interaction. This is conceptually supported by our framework, but will require an extension of the mechanisms we use for event triggering. In this case, approaches to further automate the latter should also be investigated.
-A Statement on ethical issues
No ethical issues were raised by this work. Our device profiling activities are in line with the EU legislation on reverse engineering [47, 48]. All experiments were performed in a controlled environment, and did not involve any user personal data or other privacy concerns.
-B Additional results
Fig. 11 shows the count of unique Flow IDs discovered by our framework for each investigated event, split into first-level and hidden Flow IDs. The x-axis pertains to the event ID, as provided in Table I.
Figure 11: Number of unique Flow IDs per event.